# Topological piezomagnetic effect in two-dimensional Dirac quadrupole altermagnets

### v1

@@ Abstract
Altermagnets are a recently proposed class of magnetic materials which combine properties of both ferromagnets (e.g., non-relativistic spin splitting) and antiferromagnets (e.g., compensated collinear magnetic order) Šmejkal et al. (2022a, b); Mazin and Editors (2022); Turek (2022); Mazin et al. (2021); Song et al. (2025); Jungwirth et al. (2025, 2026); Hayami et al. (2019); Bai et al. (2023); Yuan et al. (2021); Karube et al. (2022); Egorov and Evarestov (2021); Guo et al. (2023); Naka et al. (2019); Yuan and Zunger (2023); Hayami et al. (2020); González-Hernández et al. (2021); Naka et al. (2021); Shao et al. (2021); Ma et al. (2021); Bose et al. (2022). This unusual combination of properties is enabled by the particular crystallographic environment of the magnetic atoms, and signals that crystal structure, magnetic order, and symmetry are interrelated in unconventional yet fundamental ways in altermagnets. One direct consequence of this interrelation is that altermagnets are natural candidates for piezomagnetic responses, i.e., the magnetization response to applied strain (sketched in Fig. 1Ma et al. (2021); McClarty and Rau (2024); Aoyama and Ohgushi (2024); Steward et al. (2023); Fernandes et al. (2024); Takahashi et al. (2025); Naka et al. (2025); Khodas et al. (2025). This magnetization response generally has two contributions, the spin and the orbital contribution, and here we focus exclusively on the latter. The orbital magnetization is sensitive to the quantum geometric structure of the eigenstate wave functions, in particular the Berry curvature, and therefore carries a signature of band topology Thonhauser et al. (2005); Xiao et al. (2005); Ceresoli et al. (2006); Mitscherling et al. (2025).
@@
where $\bm{\tau}=(\tau^{x},\tau^{y},\tau^{z})$ is a set of Pauli matrices associated with the orbital degree of freedom. Here ${\varepsilon}_{\bf k}$ describes an orbital-independent dispersion given by ${\varepsilon}_{\bf k}=-t_{0}(\cos k_{x}+\cos k_{y})$ and the functions $t_{z,{\bf k}}$ and $t_{x,{\bf k}}$ are given by $t_{z,{\bf k}}=t_{1}(2-\cos k_{x}-\cos k_{y})-\delta$ and $t_{x,{\bf k}}=4t_{d}\sin k_{x}\sin k_{y}$, respectively. These terms correspond to an intra-orbital nearest neighbor hopping amplitude $(t_{0}\pm t_{1})/2$ for the $s$ and $d$ bands (with $t_{1}>t_{0}$) and a next-nearest neighbor interorbital hopping $t_{d}$. The form of $t_{x,{\bf k}}$ is imposed by the symmetry of the $d_{xy}$ orbital, such that $H_{0,{\bf k}}$ respects all symmetries of the square lattice when $\Delta=0$. The parameter $\delta$ controls the relative energy of the $s$ and $d$ bands, and determines whether the bands are inverted ($\delta>0$) or uninverted ($\delta<0$). In the inverted regime and the fully symmetric limit $\Delta=0$, $H_{0,{\bf k}}$ realizes a Dirac quadrupole semimetal with Dirac points on the $k_{x}=0$ and $k_{y}=0$ lines, as sketched in Fig. 1(b) Hirsbrunner et al. (2024). The coupling between the $s$ and $d_{xy}$ states described by $\Delta$ represents a symmetry breaking term which gaps these Dirac points. It breaks both time-reversal ($\mathcal{T}$) symmetry and all spatial symmetries under which the $d_{xy}$ orbital is odd, in particular the fourfold rotation $\mathcal{C}_{4z}$. Crucially, the product of time-reversal and fourfold rotation, denoted $\mathcal{T}\mathcal{C}_{4z}$, is preserved, and in this sense Eq. (1) describes a spinless system in the same symmetry class as a $d$-wave altermagnet Šmejkal et al. (2022a). It then follows that from a symmetry perspective this “orbital altermagnet” supports the same orbital magnetization response to applied strain expected in spinful $d$-wave altermagnets.
@@
such that strain effects are specified by the functions $\chi_{\bf k}$ and ${\bf w}_{\bf k}$, and ${\bf n}_{\bf k}=(t_{x,{\bf k}},\Delta,t_{z,{\bf k}})$. In the orbital altermagnet model it is sufficient to choose $\chi_{\bf k}=-2t_{0}(\cos k_{x}-\cos k_{y})$ and ${\bf w}_{\bf k}=0$, which corresponds to a change of the hopping $t_{0}\rightarrow t_{0}(1\pm\phi)$ along the $x/y$ direction. Additional symmetry-allowed strain couplings—which do not affect the topological contribution to the polarizability—are discussed in Appendix A. With the parametrization of Eq. (3), and given the specific choice of strain coupling (${\bf w}_{\bf k}=0$), the orbital piezomagnetic polarizability takes the form Venderbos (2025)
$\Lambda=-\frac{e}{\hbar}\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}~\chi_{\bf k}\frac{{\bf n}_{\bf k}\cdot\partial_{x}{\bf n}_{\bf k}\times\partial_{y}{\bf n}_{\bf k}}{2|{\bf n}_{\bf k}|^{3}},$ (4)
which has a straightforward interpretation: it is the integral over the Berry curvature of the occupied valence band multiplied by $\chi_{\bf k}$, the function encoding the strain coupling. As such, $\Lambda$ is sensitive to the topology of the occupied band.
@@
In Fig. 2(b) we show $\Lambda$ as a function of $\Delta$ for different values of $\delta$, computed using Eq. (4) (see also Appendix A). A striking feature observed in Fig. 2(b) is the finite limiting value of $\Lambda$ as $\Delta\rightarrow 0$ when $\delta>0$ (i.e., in the band-inverted regime), giving rise to a discontinuity at $\Delta=0$. This is counterintuitive, since $\Delta\rightarrow 0$ corresponds to a limit in which $\mathcal{T}$, which forbids a piezomagnetic response, is restored. The nonzero polarizability in the limit of vanishing altermagnetic order $\Delta$ is a direct consequence of the Dirac quadrupole semimetallic parent state of the orbital altermagnet, and a manifestation of the topological response theory governing topological semimetals in two dimensions Ramamurthy and Hughes (2015).
@@
Figure 3: (a) Lieb lattice altermagnet model. The non-magnetic sites are shown in grey and the magnetic sites are shown in red and blue (AA and $B$ sublattice, respectively), indicating opposite alignment of spin moments along the $\hat{z}$ direction. (b) Location of the Dirac points on the BZ boundary in the topological regime ($|N_{z}|<4t_{d}$). Valleys labeled red and blue correspond to the $\sigma=\uparrow$ and $\sigma=\downarrow$ sectors, respectively (assuming $N_{z}>0$, see Appendix B). (c) Effect of strain on the (tilted) Dirac points. The Dirac points of the valleys $\mathbf{K}_{\pm}$ and $\mathbf{K}^{\prime}_{\pm}$ are shifted in energy in opposite directions. (d) Orbital piezomagnetic polarizability as a function of $\lambda$ for different $N_{z}$ (both in units of $t_{1}$), calculated using Eq. (8). Here we have set $(t_{0},t_{d})=(0.5t_{1},1.5t_{1})$. (e) Same as in (d), showing contributions from the geometric and interband terms in Eq. (8) (for $N_{z}=2.0t_{1}$).
To examine the consequences of the Lieb lattice Dirac quadrupole structure, in particular for the piezomagnetic polarizability, we construct a linearized continuum model by expanding around the Dirac points. The Dirac points are located at ${\bf K}_{\pm}=(\pi,\pi\pm k_{D})$ and ${\bf K}^{\prime}_{\pm}=(\pi\pm k_{D},\pi)$, with $k_{D}\in(0,\pi)$ given by the solution to the equation $\sin(k_{D}/2)=\sqrt{|N_{z}|/4t_{d}}$, and thus set by the magnitude of $N_{z}$ (see Appendix B). An expansion around these points yields the Dirac Hamiltonians
@@
Lieb lattice piezomagnetic polarizability.—The expression for the orbital piezomagnetic polarizability applicable to the Lieb lattice model is a sum over contributions from each spin sector, which takes the form Venderbos (2025)
$\Lambda=-\frac{e}{\hbar}\sum_{\sigma}\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\bigg\{\chi_{\bf k}\frac{{\bf n}^{\sigma}_{\bf k}\cdot\partial_{x}{\bf n}^{\sigma}_{\bf k}\times\partial_{y}{\bf n}^{\sigma}_{\bf k}}{2|{\bf n}^{\sigma}_{\bf k}|^{3}}\\ +\bigg[\partial_{x}\varepsilon_{\bf k}\frac{{\bf w}_{\bf k}\cdot{\bf n}^{\sigma}_{\bf k}\times\partial_{y}{\bf n}^{\sigma}_{\bf k}}{2|{\bf n}^{\sigma}_{\bf k}|^{3}}-(x\leftrightarrow y)\bigg]\bigg\}.$ (8)
In each spin sector the polarizability is a sum of two terms, the “geometric term” and the “interband term”, the second of which appears when strain also affects the band mixing and ${\bf w}_{\bf k}$ is nonzero Venderbos (2025). Note that the geometric term is the only term contributing in Eq. (4).
@@
In Fig. 3(d) we show the piezomagnetic polarizability of the Lieb lattice model as a function of $\lambda$, the spin-orbit interaction, for different values of $N_{z}$. It exhibits a discontinuity at $\lambda=0$ similar to Fig. 2(b), which is at odds with the expectation that an orbital piezomagnetic effect is forbidden when full spin-rotation invariance is restored in the limit $\lambda\rightarrow 0$. As was the case for the orbital altermagnet, the finite limiting value of $\Lambda$ corresponds to a topological contribution to the polarizability and originates from the Dirac quadrupole parent state. The topological contribution can be obtained from a calculation of $\Lambda$ within the continuum Dirac model given by Eqs. (6) and (7), combined with the strain terms $\mathcal{W}_{{\bf K}_{\pm}}$ and $\mathcal{W}_{{\bf K}^{\prime}_{\pm}}$. As detailed in Appendix C, this yields $\Lambda\simeq\text{sgn}(\lambda N_{z})(e/\pi\hbar)\left(w^{D}_{0}+v_{0}w^{D}_{z}/v_{2}\right)=\text{sgn}(\lambda N_{z})(4t_{0}e/\pi\hbar)$, and reveals that both the geometric term and the interband term of Eq. (8) give rise to a topological contribution (proportional to $w^{D}_{0}$ and $w^{D}_{z}$, respectively). This is evidenced by Fig. 3(e), which separately shows the contributions of the geometric and interband terms to $\Lambda$. The topological contribution from the geometric term can be understood as before from the way applied strain shifts the Lieb lattice Dirac points in energy and creates a Dirac dipole. This is shown in Fig. 3(c). The contribution from the interband term is a consequence of the interpretation of the interband term in terms of a generalized band curvature, which is specific to two-band systems. Furthermore, in two-band Dirac models the generalized curvature is proportional to the standard Berry curvature, and yields a topological contribution to $\Lambda$ when the Dirac cones are tilted, as sketched Fig. 3(c) (see Appendix C).
@@
Discussion and conclusion.—The central result of this paper is the prediction of a topological orbital piezomagnetic effect in 2D altermagnets described by the topological response theory of Dirac semimetals. According to this response theory, the orbital magnetization $M_{z}$ of a Dirac semimetals in 2D is proportional to the time component of the Dirac point energy-momentum dipole moment. Since the latter is generically not quantized, the response theory of topological semimetals describes non-quantized responses, in contrast to the quantized responses of topological insulators.
@@
Here we have examined two minimal lattice models of Dirac quadrupole altermagnets. The two-band orbital altermagnet model, in which the Dirac quadrupole arises as a result of a band inversion, clearly exposes the orbital nature of the effect—since spin is absent—and provides the simplest nontrivial model for studying the interplay of altermagnetism and topology. The Lieb lattice model has not only garnered broad recognition as a paradigmatic minimal model for studying a variety of different phenomena related to altermagnetism, but also establishes a direct connection to materials of active current interest. A number of materials with Lieb lattice motif have been identified or proposed Chang et al. (2025), providing a realistic platform for experimentally probing orbital piezomagnetism. A promising family of materials includes the layered bulk compounds V2Se2O and V2Te2O, as well their Rb or K-intercalated variants Lin et al. (2018); Ma et al. (2021); Li et al. (2024); Jiang et al. (2025); Zhang et al. (2025), which realize the Lieb lattice in monolayer form. The correlated insulator La2O3Mn2Se2 is another confirmed altermagnet in which the MnO2 planes form a Lieb lattice Wei et al. (2025). A promising proposed metallic compound is Sr2CrO2Cr2OAs2, which is predicted to exhibit a large altermagentic magnon splitting Chang et al. (2025).
@@
Appendix A Details of the orbital altermagnet model
This Appendix collects further details on the analysis of the orbital altermagnet model. First, we present a derivation of the continuum Dirac quadrupole model. Next we discuss a more general version of the strain coupling, and the conclude with explicit formulas for the piezomagnetic polarizability $\Lambda$.
Continuum Dirac quadrupole model. The first step in the derivation of a continuum theory is to determine the location of the Dirac points in the limit $\Delta=0$. The momenta where the Dirac points are located are the solutions of the equation $t_{z,{\bf k}}=t_{1}(2-\cos k_{x}-\cos k_{y})-\delta=0$, subject to the constraint $k_{x}=0$ or $k_{y}=0$. Consider first the case $k_{x}=0$. The equation we seek to solve reduces to $t_{1}(1-\cos k_{y})-\delta=2t_{1}\sin^{2}(k_{y}/2)-\delta=0$, from which it directly follows that

### v2

@@ Abstract
Altermagnets are a recently proposed class of magnetic materials which combine properties of both ferromagnets (e.g., non-relativistic spin splitting) and antiferromagnets (e.g., compensated collinear magnetic order) Šmejkal et al. (2022a, b); Mazin and Editors (2022); Turek (2022); Mazin et al. (2021); Song et al. (2025); Jungwirth et al. (2025, 2026); Hayami et al. (2019); Bai et al. (2023); Yuan et al. (2021); Karube et al. (2022); Egorov and Evarestov (2021); Guo et al. (2023); Naka et al. (2019); Yuan and Zunger (2023); Hayami et al. (2020); González-Hernández et al. (2021); Naka et al. (2021); Shao et al. (2021); Ma et al. (2021); Bose et al. (2022). This unusual combination of properties is enabled by the particular crystallographic environment of the magnetic atoms, and signals that crystal structure, magnetic order, and symmetry are interrelated in unconventional yet fundamental ways in altermagnets. One direct consequence of this interrelation is that altermagnets are natural candidates for piezomagnetic responses, i.e., the magnetization response to applied strain (sketched in Fig. 1Ma et al. (2021); McClarty and Rau (2024); Aoyama and Ohgushi (2024); Steward et al. (2023); Fernandes et al. (2024); Takahashi et al. (2025); Naka et al. (2025); Khodas et al. (2025). This magnetization response generally has two contributions, the spin and the orbital contribution 1, and here we focus exclusively on the latter. The orbital magnetization is sensitive to the quantum geometric structure of the eigenstate wave functions, in particular the Berry curvature, and therefore carries a signature of band topology Thonhauser et al. (2005); Xiao et al. (2005); Ceresoli et al. (2006); Mitscherling et al. (2025).
@@
where $\bm{\tau}=(\tau^{x},\tau^{y},\tau^{z})$ is a set of Pauli matrices associated with the orbital degree of freedom. Here ${\varepsilon}_{\bf k}$ describes an orbital-independent dispersion given by ${\varepsilon}_{\bf k}=-t_{0}(\cos k_{x}+\cos k_{y})$ and the functions $t_{z,{\bf k}}$ and $t_{x,{\bf k}}$ are given by $t_{z,{\bf k}}=t_{1}(2-\cos k_{x}-\cos k_{y})-\delta$ and $t_{x,{\bf k}}=4t_{d}\sin k_{x}\sin k_{y}$, respectively. These terms correspond to an intra-orbital nearest neighbor hopping amplitude $(t_{0}\pm t_{1})/2$ for the $s$ and $d$ bands (with $t_{1}>t_{0}$) and a next-nearest neighbor interorbital hopping $t_{d}$. The form of $t_{x,{\bf k}}$ is imposed by the symmetry of the $d_{xy}$ orbital, such that $H_{0,{\bf k}}$ respects all symmetries of the square lattice when $\Delta=0$. The parameter $\delta$ controls the relative energy of the $s$ and $d$ bands, and determines whether the bands are inverted ($\delta>0$) or uninverted ($\delta<0$). In the inverted regime and the fully symmetric limit $\Delta=0$, $H_{0,{\bf k}}$ realizes a Dirac quadrupole semimetal with Dirac points on the $k_{x}=0$ and $k_{y}=0$ lines, as sketched in Fig. 1(b) Hirsbrunner et al. (2024). The coupling between the $s$ and $d_{xy}$ states described by $\Delta$ represents a symmetry-breaking term which gaps these Dirac points. It breaks both time-reversal ($\mathcal{T}$) symmetry and all spatial symmetries under which the $d_{xy}$ orbital is odd, in particular the fourfold rotation $\mathcal{C}_{4z}$. Crucially, the product of time-reversal and fourfold rotation, denoted $\mathcal{T}\mathcal{C}_{4z}$, is preserved, and in this sense Eq. (1) describes a spinless system in the same symmetry class as a $d$-wave altermagnet Šmejkal et al. (2022a); Chakraborty et al. (2025); Pan et al. (2025). It then follows that from a symmetry perspective this “orbital altermagnet” supports the same orbital magnetization response to applied strain expected in spinful $d$-wave altermagnets.
@@
such that strain effects, in general, are specified by the functions $\chi_{\bf k}$ and ${\bf w}_{\bf k}=(w_{x,{\bf k}},w_{y,{\bf k}},w_{z,{\bf k}})$  [^1] Basic symmetry requirements suggest that $w_{y,{\bf k}}\equiv 0$., and ${\bf n}_{\bf k}=(t_{x,{\bf k}},\Delta,t_{z,{\bf k}})$. With this parametrization of the Hamiltonian and the strain perturbation the orbital piezomagnetic polarizability takes the form Venderbos (2025)
The polarizability is a sum of two terms, the “geometric term” and the “interband term” Venderbos (2025), and the former has a straightforward interpretation: it is the integral over the Berry curvature of the occupied valence band multiplied by $\chi_{\bf k}$, the function encoding the strain coupling. This signals that $\Lambda$ is sensitive to the topology of the occupied band. To compute $\Lambda$ for the orbital altermagnet model, we choose a strain coupling given by $\chi_{\bf k}=-t_{0}(\cos k_{x}-\cos k_{y})$ and ${\bf w}_{\bf k}=0$. This corresponds to a change of the hopping $t_{0}\rightarrow t_{0}(1\pm\phi)$ along the $x/y$ direction, and since ${\bf w}_{\bf k}=0$, only the first term (i.e., the ‘geometric” term) in (4) contributes. More general strain couplings are discussed in Appendix B.
@@
In Fig. 2(b) we show $\Lambda$ as a function of $\Delta$ for different values of $\delta$. A striking feature observed in Fig. 2(b) is the finite limiting value of $\Lambda$ as $\Delta\rightarrow 0$ when $\delta>0$ (i.e., in the band-inverted regime), giving rise to a discontinuity at $\Delta=0$. This is counterintuitive, since $\Delta\rightarrow 0$ corresponds to a limit in which $\mathcal{T}$, which forbids a piezomagnetic response, is restored. The nonzero polarizability in the limit of vanishing altermagnetic order $\Delta$ is a direct consequence of the Dirac quadrupole semimetallic parent state of the orbital altermagnet, and a manifestation of the topological response theory governing topological semimetals in two dimensions Ramamurthy and Hughes (2015).
@@
Figure 3: (a) Lieb lattice altermagnet model. The non-magnetic sites are shown in grey and the magnetic sites are shown in red and blue (AA and $B$ sublattice, respectively), indicating opposite alignment of spin moments along the $\hat{z}$ direction. (b) Location of the Dirac points on the BZ boundary in the topological regime ($|N_{z}|<4t_{d}$). Valleys labeled red and blue correspond to the $\sigma=\uparrow$ and $\sigma=\downarrow$ sectors, respectively (assuming $N_{z}>0$, see Appendix D). (c) Effect of strain on the (tilted) Dirac points. The Dirac points of the valleys $\mathbf{K}_{\pm}$ and $\mathbf{K}^{\prime}_{\pm}$ are shifted in energy in opposite directions. (d) Orbital piezomagnetic polarizability as a function of $\lambda$ for different $N_{z}$ (both in units of $t_{1}$), calculated using a generalization of Eq. (4) (see Appendix E). Here we have set $(t_{0},t_{d})=(0.5t_{1},1.5t_{1})$. (e) Same as in (d), showing contributions from the geometric and interband terms (for $N_{z}=2.0t_{1}$).
To examine the consequences of the Lieb lattice Dirac quadrupole structure, in particular for the piezomagnetic polarizability, we construct a linearized continuum model by expanding around the Dirac points. The Dirac points are located at ${\bf K}_{\pm}=(\pi,\pi\pm k_{D})$ and ${\bf K}^{\prime}_{\pm}=(\pi\pm k_{D},\pi)$, with $k_{D}\in(0,\pi)$ given by the solution to the equation $\sin(k_{D}/2)=\sqrt{|N_{z}|/4t_{d}}$, and thus set by the magnitude of $N_{z}$ (see Appendix D). An expansion around these points yields the Dirac Hamiltonians
@@
Lieb lattice piezomagnetic polarizability.—The expression for the orbital piezomagnetic polarizability applicable to the Lieb lattice model is a sum over contributions from each spin sector, with the contribution from each spin sector given by the two-band formula of Eq. (4) (see Appendix E). In Fig. 3(d) we show the piezomagnetic polarizability of the Lieb lattice model as a function of $\lambda$, the spin-orbit interaction, for different values of $N_{z}$. It exhibits a discontinuity at $\lambda=0$ similar to Fig. 2(b), which is at odds with the expectation that an orbital piezomagnetic effect is forbidden when full spin-rotation invariance is restored in the limit $\lambda\rightarrow 0$. As was the case for the orbital altermagnet, the finite limiting value of $\Lambda$ corresponds to a topological contribution to the polarizability and originates from the Dirac quadrupole parent state. The topological contribution can be obtained from a calculation of $\Lambda$ within the continuum Dirac model given by Eqs. (6) and (7), combined with the strain terms $\mathcal{W}_{{\bf K}_{\pm}}$ and $\mathcal{W}_{{\bf K}^{\prime}_{\pm}}$. As detailed in Appendix E, this yields $\Lambda\simeq\text{sgn}(\lambda N_{z})(e/\pi\hbar)\left(w^{D}_{0}+v_{0}w^{D}_{z}/v_{2}\right)=\text{sgn}(\lambda N_{z})(4t_{0}e/\pi\hbar)$, and reveals that both the geometric term and the interband term give rise to a topological contribution (proportional to $w^{D}_{0}$ and $w^{D}_{z}$, respectively). This is evidenced by Fig. 3(e), which separately shows the contributions of the geometric and interband terms to $\Lambda$. The topological contribution from the geometric term can be understood as before from the way applied strain shifts the Lieb lattice Dirac points in energy and creates a Dirac dipole. This is shown in Fig. 3(c). The contribution from the interband term is a consequence of its interpretation in terms of a generalized band curvature Venderbos (2025), which is specific to two-band systems. In the case of two-band Dirac models this generalized curvature is proportional to the standard Berry curvature, and yields a topological contribution to $\Lambda$ provided the Dirac cones are tilted, as sketched in Fig. 3(c) (see Appendix E).
@@
Discussion and conclusion.—The central result of this paper is the prediction of a topological orbital piezomagnetic effect in 2D altermagnets described by the topological response theory of Dirac semimetals. According to this response theory, the orbital magnetization $M_{z}$ of a Dirac semimetal in 2D is proportional to the time component of the Dirac point energy-momentum dipole moment. Since the latter is generically not quantized, the response theory of topological semimetals describes non-quantized responses, in contrast to the quantized responses of topological insulators.
@@
Here we have examined two minimal lattice models of Dirac quadrupole altermagnets. The two-band orbital altermagnet model, in which the Dirac quadrupole arises as a result of a band inversion, clearly exposes the orbital nature of the effect—since spin is absent—and provides the simplest nontrivial model for studying the interplay of altermagnetism and topology. The Lieb lattice model has not only garnered broad recognition as a paradigmatic minimal model for studying a variety of different phenomena related to altermagnetism, but also establishes a direct connection to materials of active current interest. A number of materials with Lieb lattice motif have been identified or proposed Chang et al. (2025), providing a realistic platform for experimentally probing orbital piezomagnetism. A promising family of materials includes the layered bulk compounds V2Se2O and V2Te2O, as well their Rb or K-intercalated variants Lin et al. (2018); Ma et al. (2021); Li et al. (2024); Jiang et al. (2025); Zhang et al. (2025), which realize the Lieb lattice in monolayer form. The correlated insulator La2O3Mn2Se2 is another confirmed altermagnet in which the MnO2 planes form a Lieb lattice Wei et al. (2025). A promising proposed metallic compound is Sr2CrO2Cr2OAs2, which is predicted to exhibit a large altermagnetic magnon splitting Chang et al. (2025).
@@
Appendix A Linearized continuum model for the orbital altermagnet
This Appendix collects further details on the construction of a continuum Dirac model for the orbital altermagnet. First, we perform an expansion of the lattice model around the Dirac points in the absence of strain. Then we determine how the strain term considered in the main text enters the continuum Dirac model. In Appendix B we develop a more general theory of strain coupling.
Construction of continuum Dirac model without strain. The first step in the derivation of a continuum theory is to determine the location of the Dirac points in the limit $\Delta=0$. The momenta where the Dirac points are located are the solutions of the equation $t_{z,{\bf k}}=t_{1}(2-\cos k_{x}-\cos k_{y})-\delta=0$, subject to the constraint $k_{x}=0$ or $k_{y}=0$. Consider first the case $k_{x}=0$. The equation we seek to solve reduces to $t_{1}(1-\cos k_{y})-\delta=2t_{1}\sin^{2}(k_{y}/2)-\delta=0$, from which it directly follows that